A 5.60 gram sample is heated to an initial temperature of 100 oC. The sample is then placed in a calorimeter containing 71 grams of water starting at 24 oC. If the final temperature is 27 oC what is the specific heat of the sample?

sH2O = 4.18


1
Expert's answer
2021-02-23T04:27:10-0500

Q= C x m x deltaT

sH2O = 4.18 J/(kg °C)

Q(water) = 4.18 x 0.071 x (27-24) = 0.9 J


C = Q/(m x deltaT) = 0.9/(0.0056 x (100-27)) = 2.2 J/(kg °C)